OpenStudy (anonymous):

The sum of two numbers is 145. Their difference is 45. What are the numbers?

OpenStudy (campbell_st):

x + y = 145 x - y = 45 add the equations gives 2x = 190 x = 95 use this to find y

OpenStudy (anonymous):

would the answer be 50 and 95?/

OpenStudy (campbell_st):

they add to 145

OpenStudy (anonymous):

yes they do

OpenStudy (campbell_st):

then there is the answer

OpenStudy (anonymous):

thanx alot

OpenStudy (anonymous):

How many solutions does the system of equations below have? 2y - x = 5 3x - 6y = 5

OpenStudy (campbell_st):

in... they are both straight lines so will only intersect at one point... so the solution is an ordered pair